OpenStudy (anonymous):

Amanda earned a score of 940 on a national achievement test that was normally distributed. The mean test score was 850 with a standard deviation of 100. If 1000 students took the test, how many scored below Amanda? Use your z table. Question 6 options: 320 180 100 820

OpenStudy (anonymous):

@graciek

OpenStudy (graciek):

The z-score for 940 is found by using this formula z=X−μσ=940−850100=you can calculate When you have the z-score, use the table to find the cumulative probability. Then multiply the cumulative proability by 1000 to find how many scored below Amanda.?

OpenStudy (graciek):

i honestly just looked it up love

OpenStudy (anonymous):

i dont really know should i pick d

OpenStudy (graciek):

i think so? im not really sure :/

OpenStudy (anonymous):

Use the z-score formulae to get the value \[z = \frac{(X - \mu)}{\sigma}\] Use this and remember one thing that it was asking how mnay scored less than the her score not less than equal to.

OpenStudy (anonymous):

@Britbrat1997

OpenStudy (britbrat1997):

z=(940−850)/1000 im not positive that, that is correct it its my best guess as to what the formula would be

OpenStudy (anonymous):

so its 820

OpenStudy (britbrat1997):

The first step to solving this question is to figure out the difference between Amanda's score and the mean score: 940 minus 850 is 90. This tells us how different her score was from the average. Since the standard deviation is 100, we know that Vivian's difference (90) is within one standard deviation from the mean. In other words, if you increase the mean by one standard deviation, 850 + 100, you get 950, and you can see that Vivian's score of 940 doesn't get that high. Specifically, her difference of 90 is 90% of the standard deviation of 100. The decimal equivalent of 90% is 0.90, and this is called Vivian's z-score.In math notation, we've done this: z = (X - μ) / σ = (940 - 850) / 100 = 0.90 In math notation, we've done this: z = (X - μ) / σ = (940 - 850) / 100 = 0.90. As you may know, in a normal distribution it's expected that about 68% of all observations will fall within 1 standard deviation of the mean, 95% will fall within 2 standard deviations, and 99% will fall within 3 standard deviations. You now have all the info you need to do a rough calculation of the cumulative probability. Here's how to put it together: The upper half of the 68% (the scores between the mean and the first standard deviation) = 34% Vivian's 90% of the 34% = 30.6% All the test scores lower than the mean = 0.500 (this is 1/2 of all test scores) + All the test scores between the mean and Vivian's score = 0.306 (see above) = 0.806, which is not the real cumulative probability, but an approximation. source; yahoo answers

OpenStudy (anonymous):

Mine answer does not match to anyone. As its asking for score less than Amanda so we have to take less random distribution variable. And the answer will change after that . Its coming close to 813 but no answer has 813 value.

OpenStudy (anonymous):

so the best response would b b right

OpenStudy (anonymous):

no wait its b right @baljeet16

OpenStudy (anonymous):

i meant @Britbrat1997

OpenStudy (britbrat1997):

Vivian's 90% of the 34% = 30.6% All the test scores lower than the mean = 0.500 (this is 1/2 of all test scores) + All the test scores between the mean and Vivian's score = 0.306 (see above) = 0.806, which is not the real cumulative probability, but an approximation. My guess would be the answer that is in the 800, only because of this answer off of yahoo. Personally i could not find an answer that matched any of the given choices above

OpenStudy (anonymous):

I think either data is incorrect or options are.

OpenStudy (britbrat1997):

it could be, you should just be able to plug it into the equation and check it to see if you get the correct answer